# On misconceptions about the Brier score in binary prediction models

### v3

@@ Appendix A Why we care about the expectation: law of large numbers and central limit theorem
Furthermore, the individual terms $(Q_{i}-Y_{i})^{2}$ have finite variance and hence by the CLT the distribution of the BS, when properly normalised, approaches a normal distribution.
@@
Appendix B More detail for mathematical understanding BS
B.1. Expectation of Brier score in one dimension
@@
B.2. Differences in expectation of Brier score in one dimension
@@
B.3. Understanding the expectation of the general BS
@@
(B.2) $f((1-\alpha)x+\alpha y)\geq(1-\alpha)f(x)+\alpha f(y)$
Now we come back to the case where we have $n$ observations and we consider the BS.
Lemma B.1.
@@
Appendix C Mathematical proof impossibility of Brier score 0
@@
Appendix D MAE and CIL are not strictly proper
Recall the formulas from (4.2) and (4.3), and let $>\sim Bern(q),q\in[0,1]$. It is enough to do the case $n=1$ for MAE and CIL, the case for more obsevations follows analogous. We have to show that the expectation is not uniquely minimised (resp maximised if the orientation is opposite) for $p\in[0,1]$ if we choose $p=q$ [20].
  • We calculate
    $\mathbb{E}[MAE(p,Y)]=qp+(1-q)(1-p),$
    to simplify further let $q=1/2$, then $\mathbb{E}[MAE(p,Y)]=1/2,$ and the expectation does not depend on $p$.
  • We calculate
    $\mathbb{E}[CIL(p,Y)]=p-q$
    which is clear.
Appendix E More details for the simulation via the ADEMP framework
@@
E.1. Aims
  • The main aim is to compare Brier score across different settings both for true probabilities as well as predictions, which are based on the perfect prediction with some noise or bias added through median.
  • A secondary aim is to compare Brier score to the mean incident BS, to compare it to CIL. In particular to estimate the probability that $\bar{y}-\bar{y}^{2}>BS_{perf}$, and to estimate median of $\bar{y}-\bar{y}^{2}-BS_{perf}$.
E.2. Data Generating Mechanism
E.2.1. $y_{i}$ entered in the BS
@@
  • $Unif(a,b)$, where $(a,b)\in\{(0,1),(0.3,0.7),(0,0.2)\}$.
  • $Beta(\alpha,\beta)$, where $(\alpha,\beta)\in\{(2,2),(5,5)\}$.
@@
E.2.2. Predictions $p_{i}$ entered in the BS
@@
  • $+Unif(-0.2,0.2)$: $p_{i}=q_{i}+X_{i}$, where $X_{i}\sim Unif(-0.2,0.2)$, i.e., fairly disturbed but unbiased predictions.
@@ E.2.3. Sample Size
E.2.4. Number of DGM Scenarios and Simulation Runs
@@
E.3. Estimand/Target of Analysis
  • Mean of BS.
  • 2.5% and 97.5% quantile of BS.
E.4. Methods
E.4.1. Basis of Simulations
The simulation is run as a Monte-Carlo simulation, where for the two settings based on NHANES data, the $q_{i}$ are based on the predicted value of the logistic regression for the corresponding observation, and the $q_{i}$ are subsampled without replacement.
E.5. Performance Measures
  • Median of Brier score calculated as sample median.
  • 2.5% and 97.5% quantile of Brier score calculated as corresponding quantiles.
  • CIL and its comparison to the BS.
@@ E.5. Performance Measures
  • estimate the median of $\bar{y}-\bar{y}^{2}-BS_{perf}$.

### v4

@@ Appendix B Why we care about the expectation: law of large numbers and central limit theorem
Furthermore, the individual terms $(Q_{i}-Y_{i})^{2}$ have finite variance and hence by the CLT the distribution of the Brier score, when properly normalised, approaches a normal distribution.
@@
Appendix C Connections to some other scores
We mention three other scores connected to Brier score. The Brier score from (1.1) equals mathematically the MSE, hence its square root is the root mean squared error (RMSE):
(C.1) $RMSE(p,y):=\sqrt{BS(p,y)}.$
(C.2) $MAE(p,y)=\frac{1}{n}\sum_{i=1}^{n}|p_{i}-y_{i}|,$
and the CIL [35]
(C.3) $CIL(p,y)=\frac{1}{n}\sum_{i=1}^{n}p_{i}-y_{i}.$
(C.4) $CIL(p,y)\leq MAE(p,y)\leq RMSE(p,y)$
as well as
(C.5) $MSE(p,y)\leq MAE(p,y)\leq RMSE(p,y).$
Appendix D More detail for mathematical understanding Brier score
D.1. Expectation of Brier score in one dimension
@@
D.2. Differences in expectation of Brier score in one dimension
@@
D.3. Understanding the expectation of the general Brier score
@@
(D.2) $f((1-\alpha)x+\alpha y)\geq(1-\alpha)f(x)+\alpha f(y)$
Now we come back to the case where we have $n$ observations and we consider the Brier score.
Lemma D.1.
@@
Appendix E Mathematical proof impossibility of Brier score 0
@@
Appendix F More details for the simulation via the ADEMP framework
@@
F.1. Aims
  • The main aim is to compare Brier score across different settings both for true probabilities as well as predictions, which are based on the perfect prediction with some noise or bias added through median, quantile, and violin plots.
  • A secondary aim is to compare Brier score to the mean incident Brier score, as well as to compare the CIL. In particular to estimate the probability that $\bar{y}-\bar{y}^{2}>BS_{perf}$, and to visualise the observed distribution of $\bar{y}-\bar{y}^{2}-BS_{perf}$ in different settings.
F.2. Data Generating Mechanism
F.2.1. $y_{i}$ entered in the Brier score
@@
  • $Unif(a,b)$, where $(a,b)\in\{(0,1),(0,0.2)\}$.
  • $Beta(\alpha,\beta)$, where $(\alpha,\beta)\in\{(2,5),(5,5),(3,3)\}$.
@@
F.2.2. Predictions $p_{i}$ entered in the Brier score
@@
  • $+Unif(-0.1,0.1)$ resp $+Unif(-0.05,0.05)$: $p_{i}=q_{i}+X_{i}$, where $X_{i}\sim Unif(-0.1,0.1)$ resp $Unif(-0.05,0.05)$, i.e., disturbed but unbiased predictions.
@@ F.2.3. Sample Size
F.2.4. Number of DGM Scenarios and Simulation Runs
@@
F.3. Estimand/Target of Analysis
  • Distribution, median, quantiles of observed Brier score.
F.4. Methods
F.4.1. Basis of Simulations
The simulation is run as a Monte-Carlo simulation in R, where for the two settings based on NHANES data, the $q_{i}$ are based on the predicted value of the logistic regression for the corresponding observation, and the $q_{i}$ are subsampled without replacement. For the other distributions the $q_{i}$ are sampled iid from the distribution.
F.5. Performance Measures
  • distribution, median, 5% and 95% quantile Brier score in violin plot.
@@ F.5. Performance Measures
  • violin plot of $\bar{y}-\bar{y}^{2}-BS_{perf}$.